OpenStudy (anonymous):

Find k 2+3y=-1 x+4y=2 5x+ky=-4

OpenStudy (anonymous):

Okay, so, this one you'll want to start by first finding y and x. If you have two unknown variables in an equation, you can't solve for either of them. However, looking at your first equation, you should be able to begin by solving for y.

OpenStudy (anonymous):

2+3y=-1 // solve the above equation first for y. 3y = -1-2 3y = -3 y = -1 // plug that answer into the next equation to solve for x. x+4y=2 // use those two answers. plug them into the next equation and solve for k. 5x+ky=-4

OpenStudy (phi):

Is that a typo for the 1st equation?